OpenStudy (anonymous):

2log base b of x = 3 log base b of (1-a) + 2 log base b of (1+a) - log base b of (1/a -a)^2

OpenStudy (anonymous):

\[2\log_b{x}=3\log_{b}(1-a)+2\log_{b}(1+a)-\log_{b}(\frac{1}{a}-a)^2\] \[2\log_{b}x=\log_{b}(1-a)^3+\log_{b}(1+a)^2-\log_{b}(\frac{1-a^2}{a})^2\] \[2\log_{b}x=\log_{b}\frac{(1-a)^3(1+a)^2}{(\frac{1-a^2}{a})^2}\] \[2\log_{b}x=\log_{b}(\frac{a^2(1-a)^3(1+a)^2}{(1-a^2)^2})\] \[2\log_{b}x=\log_{b}(\frac{a^2(1-a^2)^2(1-a)}{(1-a^2)^2})\] \[2\log_{b}x=\log_{b}[a^2(1-a)]\] \[\log_{b}x=\frac{\log_{b}[a^2(1-a)]}{2}\] \[\large b^\frac{\log_{b}[a^2(1-a)]}{2}=x\] \[x=a^2(1-a)\]

OpenStudy (anonymous):

think i messed up on the very last line

OpenStudy (anonymous):

\[\large b^{{\log_{b}[a^2(1-a)]}^{\frac{1}{2}}}=x\] \[x=\sqrt{[a^2(1-a)]}\]

OpenStudy (anonymous):

yea...that was a lot of work...not gonna do it again...please make sure you post the question right